OpenStudy (anonymous):

Jennifer bought a brand new sports car.The price of the car was $14,900. The dealer gave her a $1,250 instant rebate on the price and 0% interest for 60 months. Jennifer made a down payment of $750. What would be the expression for Jennifer's monthly payment if shetakes 60 months to pay the remaining balances?

OpenStudy (anonymous):

14900-1250-750=12900/60=215 payed each month

OpenStudy (anonymous):

(14900 + 750 + 1250) / 60 14900 / (750 + 1250 * 60) (14900 - (750 + 1250)) / 60 (14900 - 750 - 1250) * 60 14900 + (750 - 1250) / 60 these are the choices i can choose as an answer..